# Consistency of Neural Causal Partial Identification

### v1

@@ 3 Approximation Error of Neural Causal Models
In this section, we study the expressive power of NCMs. In particular, given an SCM $\mathcal{M}^{*}$, we want to construct an NCM $\hat{\mathcal{M}}$ such that the two causal models produce similar interventional results. Unlike in the discrete case [4, 50], latent distributions can be extremely complicated in general cases. The main challenge is how to design the structure of NCMs to ensure strong approximation power.
@@
Definition 5 (Canonical representation).
@@ A SCM $\mathcal{M}$ with causal graph $\mathcal{G}$ has canonical form if
    $\boldsymbol{U}=\left\{U_{C}:C\ \text{is a}\ C^{2}\text{-component of}\ \mathcal{G}\right\}\cup\left\{G_{V_{i}}=(g^{V_{i}}_{1},\cdots,g^{V_{i}}_{n_{i}}):V_{i}\ \text{is categorical}\right\},$
@@
    $V_{i}=\begin{cases}f_{i}\left(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g^{V_{i}}_{k}+\log\left(f_{i}\left(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\|f_{i}\|_{1}=1,\end{cases}$ (3)
    where $\boldsymbol{U}_{V_{i}}=\{U_{C}:V_{i}\in C,C\ \text{is a}\ C^{2}\text{-component of}\ \mathcal{G}\}$ and $(x)_{k}$ is the $k$-th coordinate of the vector $x$. We further assume that $f_{i}$ are normalized for categorical variables.
Given a function class $\mathcal{F}$, the SCM class $\mathcal{M}(\mathcal{G},\mathcal{F},\boldsymbol{U})$ consists of all canonical SCM models with causal graph $\mathcal{G}$ such that $f_{i}\in\mathcal{F},i\in[n_{V}]$.
Proposition 4 in the appendix shows that any SCM satisfying 1 and 2 can be represented in this way. Therefore, we restrict our attention to the class $\mathcal{M}(\mathcal{G},\mathcal{F},\boldsymbol{U})$. For two SCM classes $\mathcal{M}(\mathcal{G},\mathcal{F},\boldsymbol{U}),\mathcal{M}(\mathcal{G},\hat{\mathcal{F}},\boldsymbol{\hat{U}})$, we want to study how well we can represent the models in the first class by the second class. We measure the quality of the approximation by the Wasserstein distance between the intervention distributions of the two models. To approximate the functions in the structural equations, we need to make the following regularity assumptions on the functions.
Assumption 3.
If $V_{i}$ is continuous, $f_{i}$ in (3) are $L_{f}$-Lipschitz continuous. If $V_{i}$ is categorical, the propensity functions $f_{i}(\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}})\triangleq\mathbb{P}(V_{i}=j|\text{Pa}(V_{i}),\boldsymbol{U}_{V_{i}}),j\in[n_{i}]$ are $L_{f}$-Lipschitz continuous. There exists a constant $K>0$ such that $\max_{i\in[n_{V}],j\in[n_{U}]}\{\|V_{i}\|_{\infty},\|U_{j}\|_{\infty}\}\leqslant K$.
@@
Given any SCM model $\mathcal{M}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\boldsymbol{U})$, let the treatment variable set be $\boldsymbol{T}=\{T_{k}\}_{k=1}^{n_{t}}$ and suppose that 1, 2 and 3 hold for $\mathcal{M}$ with Lipschitz constant $L$, constant $K$. For any intervention $\boldsymbol{T}=\boldsymbol{t}$ and ^$\hat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},\hat{\mathcal{F}},\hat{\boldsymbol{U}})$ , we have
$W\left(P^{\mathcal{M}}(\boldsymbol{V}(\boldsymbol{t})),P^{\hat{\mathcal{M}}}(\hat{\boldsymbol{V}}(\boldsymbol{t}))\right)\leqslant C_{\mathcal{G}}(L,K)\left(\sum_{i=1}^{n_{V}}\|f_{i}-\hat{f}_{i}\|_{\infty}+W\left(P^{\mathcal{M}}(\boldsymbol{U}),P^{\hat{\mathcal{M}}}(\hat{\boldsymbol{U}})\right)\right),$ (4)
@@
where $C_{\mathcal{G}}(L,K)$ is a constant that only depends on $L,K$ and the causal graph $\mathcal{G}$.
@@
Theorem 1 separates the approximation error into two parts, which motivates us to construct the NCM in the following way. First, we approximate the functions $f_{i}$ in (3) by NNs $\hat{f}_{i}$. Then, we approximate the distribution of latent variables by pushing forward uniform variables using neural networks, i.e., $\hat{U}_{C_{j}}=\hat{g}_{j}(Z_{C_{j}})$, where $\{C_{j}\}$ are $C^{2}$ components and $Z_{C_{j}}$ are multi-variate uniform random variables. Putting things together, the structural equations of the approximated model $\hat{\mathcal{M}}$ are
@@
$\hat{V}_{i}=\begin{cases}\hat{f}_{i}\left(\text{Pa}(\hat{V}_{i}),(\hat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\boldsymbol{U}_{V_{i}}}\right),&V_{i}\text{ is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}+\log\left(\hat{f}_{i}\left(\text{Pa}(\hat{V}_{i}),(\hat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\boldsymbol{U}_{V_{i}}}\right)\right)_{k}\right\},&V_{i}\text{ is categorical},\end{cases}$ (5)
where $Z_{C_{j}}\sim U([0,1]^{N_{C,j}})$ are i.i.d. and $N_{C,j}$ are constants to be specified later.
@@
For the first part, we need to approximate the Lipschitz continuous functions. For simplicity, we assume that the domain of the functions are uniform cubes. Similar arguments hold for any bounded cubes. We denote $\mathcal{NN}_{k_{1},k_{2}}(W,L)$ to be the set of ReLu NNs with input dimension $k_{1}$, output dimension $k_{2}$, width $W$ and depth $L$. It has been shown that $\epsilon(\mathcal{NN}_{k_{1},1}(2d_{1}+10,L_{0}),\mathcal{F}_{L}([0,1]^{d_{1}},\mathbb{R}))\leqslant O(L_{0}^{-{2}/{k_{1}}})$ [52], where $\epsilon(\cdot,\cdot)$ denotes the approximation error defined in Section 2. For a vector valued function, we can use a wider NN to approximate each coordinate and get a similar rate.
@@
For the second part, we approximate each $U_{i}$ individually by pushing forward i.i.d. multivariate uniform variables $\hat{U}_{C_{i}}=\hat{g}_{i}(Z_{C_{i}})$ since the latent variables are independent by 1. To do so, we examine under what assumptions on the measure $\mathbb{P}$ over $\mathbb{R}^{n}$ we can find a NN $\hat{g}$ such that $W(\hat{g}_{\#}\lambda,\mathbb{P})$ is small, where $\lambda(\cdot)$ is the uniform measure on cube $[0,1]^{d},d\geqslant 0$.
@@ 3.1 Approximating Mixed Distributions by Wide Neural Networks
Space-filling Curve Gumbel-Softmax Lipschitz Homeomorphism
(a) Architecture of wide neural network for $4-$dimensional output. The first (yellow) part approximates the distribution on different connected components of the support using results from [42]. The second (blue) part transforms the distributions on the unit cube to the distributions on the support. The third (green) part is the Gumbel-Softmax layer. It combines the distributions on different connected components of the support together and outputs the final distribution.
$\mathbb{P}(\boldsymbol{U})$ $\mathbb{P}(\boldsymbol{Z})$
(b) This figure demonstrates the first two parts of our architecture. Each interval in the yellow box corresponds to one coordinate of input in the left figure. We first push forward uniform distributions to different cubes. Then, using 4, we adapt the shape of the support and push the measure from unit cubes to the original support of $P(\boldsymbol{U})$. In this way, we can approximate complicated measures by pushing forward uniform variables.
@@
Assumption 4 (Mixed Distribution).
@@
4 encompasses almost all natural distributions. For example, distributions supported on $[0,1]^{d}$ and closed balls, finitely supported distributions and mixtures of them all satisfy this assumption. 4 allows us to transform the support of the targeted distribution into unit cubes and the nice geometric properties of unit cubes facilitate our construction.
@@
Now, we briefly explain the construction of the push-forward maps. The NN architecture consists of three parts (see Figure 1(a)). The input dimension is the same as the number of connected components of the support $N_{C}$. For each component $C_{i}$, let $H_{i}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$, where $d_{i}^{C}$ is the dimension of component $C_{i}$ in 4. By $\mathbb{P}=(H_{i})_{\#}(H_{i}^{-1})_{\#}\mathbb{P}$ on $C_{i}$, we can approximate $(H_{i}^{-1})_{\#}\mathbb{P}$ first, which is supported on a unit cube. [42] constructs a wide NN $\hat{g}$ of width $W$ and constant depth such that $W(\hat{g}_{\#}\lambda,(H_{i})^{-1}_{\#}\mathbb{P})\leqslant O(W^{-1/d_{i}^{C}})$ where $\lambda$ is the uniform measure on $[0,1]$. Then, we approximate the Lipschitz map $H_{i}$ to $C_{i}$ to pull the distribution back to $C_{i}$. These are the first two parts (yellow and blue blocks in Figure 1(a)) of the NN.
Suppose that the output of $i$-th coordinate in the first two parts is $v_{i}$, the Gumbel-softmax layer in the third part (green box) combines different components of the support. In particular, we want to output $v_{i}$ with probability $p_{i}=\mathbb{P}(C_{i})$. Let $V=[v_{1},\cdots,v_{i}]$, this can be achieved by $V\cdot X$, where $X=(X_{1},\cdots,X_{N_{C}})^{\mathsf{T}}$ is a one-hot random vector with $\mathbb{P}({X}_{i}=1)=p_{i}$. To use backpropagation in training, we use the Gumbel-Softmax trick [30] to (approximately) simulate such a random vector,
$\hat{X}^{\tau}_{i}=\frac{\exp((\log p_{i}+G_{i})/\tau)}{\sum_{k=1}^{N_{C}}\exp((\log p_{k}+G_{k})/\tau)},$
@@
where $\tau>0$ is the temperature parameter and $G_{i}\sim\text{Gumbel}(0,1)$ are i.i.d. standard Gumbel variables. As $\tau\rightarrow 0$, the distribution of $\hat{X}^{\tau}$ converges almost surely to the categorical distribution [35, Proposition 1]. In particular, when $\tau=0$, we denote $\hat{X}_{i}^{0}=\mathbb{I}_{i=\arg\max_{j}\{\log p_{i}+G_{i}\}}$. The output of the last layer is $V\cdot\hat{X}^{\tau}$. Note that the Gumbel-softmax function is differentiable with respect to parameter $\{\log(p_{i})\}_{i=1,\cdots,N_{C}}$. Therefore, we can train the network with common gradient-based algorithms. Putting things together, we obtain the following theorem.
Theorem 2.
@@
Given any probability measure $\mathbb{P}$ on $\mathbb{R}^{d}$ that satisfies 4, let $\lambda$ be the Lebesgue measure on $[0,1]^{N_{C}}$, where $N_{C}$ in defined in 4. There exists a neural network $\hat{g}=\hat{g}_{3}^{\tau}\circ\hat{g}_{2}\circ\hat{g}_{1}$ with the above architecture (Figure 1(a)) such that
$W(\hat{g}_{\#}\lambda,\mathbb{P})\leqslant O\left(W_{1}^{-1/\max_{i}\{d_{i}^{C}\}}+L_{2}^{-2/\max_{i}\{d_{i}^{C}\}}+(\tau-\tau\log\tau)\right).$
@@
Here, $\hat{g}_{i},i=1,2$ has the separable form $\left(\hat{g}_{i}^{1}(x_{1}),\cdots,\hat{g}_{i}^{N_{C}}(x_{N_{C}})\right)$ and $\hat{g}_{1}^{j}\in\mathcal{NN}_{1,d_{j}^{C}}(W_{1},\Theta(d_{j}^{C}))$, $\ \hat{g}_{2}^{j}\in\mathcal{NN}_{d_{j}^{C},d}(\Theta(d\cdot d_{j}^{C}),L_{2}),\ j\in[N_{C}]$. $\{d_{j}^{C}\}$ are the dimension of cubes in 4. $\hat{g}_{3}^{\tau}$ is the Gumbel-Softmax layer with temperature parameter $\tau>0$.

### v2

@@ 3 Approximation Error of Neural Causal Models
In this section, we study the expressive power of NCMs, which serves as a key ingredient in proving the consistency result. In particular, given an SCM $\mathcal{M}^{*}$, we want to construct an NCM $\widehat{\mathcal{M}}$ such that the two causal models produce similar interventional results. Unlike in the discrete case [4, 51], latent distributions can be extremely complicated in general cases. The main challenge is how to design the structure of NCMs to ensure strong approximation power.
@@
Definition 3.1 (Canonical representation).
@@ A SCM $\mathcal{M}$ with causal graph $\mathcal{G}$ has canonical form if
    $\bm{U}=\left\{U_{C}:C\ \text{is a}\ C^{2}\text{-component of}\ \mathcal{G}\right\}\cup\left\{G_{V_{i}}=(g^{V_{i}}_{1},\cdots,g^{V_{i}}_{n_{i}}):V_{i}\ \text{is categorical}\right\},$
@@
    $V_{i}=\begin{cases}f_{i}\left(\text{Pa}(V_{i}),\bm{U}_{V_{i}}\right),&V_{i}\ \text{is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g^{V_{i}}_{k}+\log\left(f_{i}\left(\text{Pa}(V_{i}),\bm{U}_{V_{i}}\right)\right)_{k}\right\},&V_{i}\ \text{is categorical},\|f_{i}\|_{1}=1,\end{cases}$ (3.1)
    where $\bm{U}_{V_{i}}=\{U_{C}:V_{i}\in C,C\ \text{is a}\ C^{2}\text{-component of}\ \mathcal{G}\}$ and $(x)_{k}$ is the $k$-th coordinate of the vector $x$. We further assume that $f_{i}$ are normalized for categorical variables.
Given a function class $\mathcal{F}$, the SCM class $\mathcal{M}(\mathcal{G},\mathcal{F},\bm{U})$ consists of all canonical SCM models with causal graph $\mathcal{G}$ such that $f_{i}\in\mathcal{F},i\in[n_{V}]$.
Proposition B.1 in the appendix shows that any SCM satisfying 2.3,2.4 can be represented in this way and we provide an example in Section B.1 to illustrate how to obtain the canonical representation for a given SCM. Therefore, we restrict our attention to the class $\mathcal{M}(\mathcal{G},\mathcal{F},\bm{U})$. For two SCM classes $\mathcal{M}(\mathcal{G},\mathcal{F},\bm{U}),\mathcal{M}(\mathcal{G},\widehat{\mathcal{F}},\bm{\widehat{U}})$, we want to study how well we can represent the models in the first class by the second class. The Wasserstein distance between the intervention distributions is used to measure the quality of the approximation. To approximate the functions in the structural equations, we need to make the following regularity assumptions on the functions.
Assumption 3.2.
If $V_{i}$ is continuous, $f_{i}$ in (3.1) are $L_{f}$-Lipschitz continuous. If $V_{i}$ is categorical, the propensity functions $f_{i}(\text{Pa}(V_{i}),\bm{U}_{V_{i}})\triangleq\mathbb{P}(V_{i}=j|\text{Pa}(V_{i}),\bm{U}_{V_{i}}),j\in[n_{i}]$ are $L_{f}$-Lipschitz continuous. There exists a constant $K>0$ such that $\max_{i\in[n_{V}],j\in[n_{U}]}\{\|V_{i}\|_{\infty},\|U_{j}\|_{\infty}\}\leqslant K$.
@@
Given any SCM model $\mathcal{M}\in\mathcal{M}(\mathcal{G},\mathcal{F}_{L},\bm{U})$, let the treatment variable set be $\bm{T}=\{T_{k}\}_{k=1}^{n_{t}}$ and suppose that 2.3, 2.4 and 3.2 hold for $\mathcal{M}$ with Lipschitz constant $L$, constant $K$. For any intervention $\bm{T}=\bm{t}$ and ^$\widehat{\mathcal{M}}\in\mathcal{M}(\mathcal{G},\widehat{\mathcal{F}},\widehat{\bm{U}})$ , we have
$W\left(P^{\mathcal{M}}(\bm{V}(\bm{t})),P^{\widehat{\mathcal{M}}}(\widehat{\bm{V}}(\bm{t}))\right)\leqslant C_{\mathcal{G}}(L,K)\left(\sum_{i=1}^{n_{V}}\|f_{i}-\widehat{f}_{i}\|_{\infty}+W\left(P^{\mathcal{M}}(\bm{U}),P^{\widehat{\mathcal{M}}}(\widehat{\bm{U}})\right)\right),$ (3.2)
@@
where $C_{\mathcal{G}}(L,K)$ is a constant that only depends on $L,K$ and the causal graph $\mathcal{G}$ and $f_{i},\widehat{f}_{i}$ are structural functions of $\mathcal{M}$ and $\widehat{\mathcal{M}}$ respectively.
@@
Theorem 3.3 separates the approximation error into two parts, which motivates us to construct the NCM in the following way. First, we approximate the functions $f_{i}$ in (3.1) by NNs $\widehat{f}_{i}$. Then, we approximate the distribution of latent variables by pushing forward uniform and Gumbel variables using neural networks, i.e., $\widehat{U}_{C_{j}}=\widehat{g}_{j}(Z_{C_{j}})$, where $\{C_{j}\}$ are $C^{2}$ components and $Z_{C_{j}}$ are multi-variate uniform and Gumbel random variables. The structural equations of the resulting approximated model $\widehat{\mathcal{M}}$ are
@@
$\widehat{V}_{i}=\begin{cases}\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),(\widehat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\bm{U}_{V_{i}}}\right),&V_{i}\text{ is continuous},\\ \arg\max_{k\in[n_{i}]}\left\{g_{k}+\log\left(\widehat{f}_{i}\left(\text{Pa}(\widehat{V}_{i}),(\widehat{g}_{j}(Z_{C_{j}}))_{U_{C_{j}}\in\bm{U}_{V_{i}}}\right)\right)_{k}\right\},&V_{i}\text{ is categorical},\end{cases}$ (3.3)
wheres $N_{C,j}$ are constants to be specified later.
@@
For the first part, we need to approximate Lipschitz continuous functions. For simplicity, we assume that the domain of the functions are uniform cubes. Similar arguments hold for any bounded cubes. We denote $\mathcal{NN}_{k_{1},k_{2}}(W,L)$ to be the set of ReLu NNs with input dimension $k_{1}$, output dimension $k_{2}$, width $W$ and depth $L$. It has been shown that $\epsilon(\mathcal{NN}_{k_{1},1}(2d_{1}+10,L_{0}),\mathcal{F}_{L}([0,1]^{d_{1}},\mathbb{R}))\leqslant O(L_{0}^{-{2}/{k_{1}}})$ [53], where $\epsilon(\cdot,\cdot)$ denotes the approximation error defined in Section 2. For a vector valued function, we can use a wider NN to approximate each coordinate and get a similar rate.
@@
For the second part, we approximate each $U_{i}$ individually by pushing forward i.i.d. multivariate uniform and Gumbel variables $\widehat{U}_{C_{i}}=\widehat{g}_{i}(Z_{C_{i}})$ since the latent variables are independent by 2.3. To do so, we examine under what assumptions on the measure $\mathbb{P}$ over $\mathbb{R}^{n}$ we can find a NN $\widehat{g}$ such that $W(\widehat{g}_{\#}\lambda,\mathbb{P})$ is small, where $\lambda(\cdot)$ is some reference measure.
@@ 3.1 Approximating Mixed Distributions by Wide Neural Networks
Space-filling Curve Gumbel-Softmax Lipschitz Homeomorphism $N_{C}$= #Connected components of latent dist. Width: $W_{2}$ Depth $L_{2}$ Width: $W_{1}$ Depth $L_{1}$ Randomly output one component
Figure 1: Architecture of wide neural network for $4-$dimensional output. The first (yellow) part approximates the distribution on different connected components of the support using the results from [43]. The width and depth of each block in this part are $W_{1}$ and $L_{1}$. The second (blue) part transforms the distributions on the unit cube to the distributions on the support. The width and depth of each block in the blue part are $W_{2}$ and $L_{2}$. The third (green) part is the Gumbel-Softmax layer. It combines the distributions on different connected components of the support together and outputs the final distribution.
[…]
Figure 2: This figure demonstrates the first two parts of our architecture. Each interval in the yellow box corresponds to one coordinate of input in the left figure. We first push forward uniform distributions to different cubes. Then, using 3.4, we adapt the shape of the support and push the measure from unit cubes to the original support of $P(\bm{U})$. In this way, we can approximate complicated measures by pushing forward uniform variables.
@@
Assumption 3.4 (Mixed Distribution).
@@
3.4 encompasses almost all natural distributions. For example, distributions supported on $[0,1]^{d}$ and closed balls, finitely supported distributions and mixtures of them all satisfy this assumption. 3.4 allows us to transform the support of the targeted distribution into unit cubes and the nice geometric properties of unit cubes facilitate our construction.
@@
Now, we briefly explain the construction of the push-forward maps. An example is provided in Section B.1 to illustrate the construction. The NN architecture consists of three parts (see Figure 2). The input dimension is the same as the number of connected components of the support $N_{C}$. For each component $C_{i}$, let $H_{i}\in\mathcal{H}([0,1]^{d^{C}_{i}},C_{i})\cap\mathcal{F}_{L}([0,1]^{d^{C}_{i}},C_{i})$, where $d_{i}^{C}$ is the dimension of component $C_{i}$ in 3.4. By $\mathbb{P}=(H_{i})_{\#}(H_{i}^{-1})_{\#}\mathbb{P}$ on $C_{i}$, we can approximate $(H_{i}^{-1})_{\#}\mathbb{P}$ first, which is supported on a unit cube. [43] constructs a wide NN $\widehat{g}$ of width $W$ and constant depth such that $W(\widehat{g}_{\#}\lambda,(H_{i})^{-1}_{\#}\mathbb{P})\leqslant O(W^{-1/d_{i}^{C}})$ where $\lambda$ is the uniform measure on $[0,1]$. Then, we approximate the Lipschitz map $H_{i}$ to $C_{i}$ to pull the distribution back to $C_{i}$. These are the first two parts (yellow and blue blocks in Figure 2) of the architecture.
@@
Suppose that the output of $i$-th coordinate in the first two parts is $\vec{v}_{i}$, the Gumbel-softmax layer in the third part (green box) combines different components of the support. In particular, we want to output $\vec{v}_{i}$ with probability $p_{i}=\mathbb{P}(C_{i})$. Let $V=[\vec{v}_{1},\cdots,\vec{v}_{i}]$, this can be achieved by outputting $VX$, where $X=(X_{1},\cdots,X_{N_{C}})^{\mathsf{T}}$ is a one-hot random vector with $\mathbb{P}({X}_{i}=1)=p_{i}$. To use backpropagation in training, we use the Gumbel-Softmax trick [31] to (approximately) simulate such a random vector, $\widehat{X}^{\tau}_{i}=\frac{\exp((\log p_{i}+G_{i})/\tau)}{\sum_{k=1}^{N_{C}}\exp((\log p_{k}+G_{k})/\tau)},$ where $\tau>0$ is the temperature parameter (a hyperparameter) and $G_{i}\sim\text{Gumbel}(0,1)$ are i.i.d. standard Gumbel variables. As $\tau\rightarrow 0$, the distribution of $\widehat{X}^{\tau}$ converges almost surely to the categorical distribution [36, Proposition 1]. In particular, when $\tau=0$, we denote $\widehat{X}_{i}^{0}=\mathbb{I}_{i=\arg\max_{j}\{\log p_{i}+G_{i}\}}$. The output of the last layer is $V\widehat{X}^{\tau}$. Note that the Gumbel-softmax function is differentiable with respect to parameter $\{\log(p_{i})\}_{i=1,\cdots,N_{C}}$. Therefore, we can train the network with common gradient-based algorithms. Putting things together, we obtain the following theorem.
Theorem 3.5.
@@
Given any probability measure $\mathbb{P}$ on $\mathbb{R}^{d}$ that satisfies 3.4, let $\lambda$ be the Lebesgue measure on $[0,1]^{N_{C}}$, where $N_{C}$ is defined in 3.4. There exists a neural network $\widehat{g}=\widehat{g}_{3}^{\tau}\circ\widehat{g}_{2}\circ\widehat{g}_{1}$ with the above architectsure (Figure 2) such that
$W(\widehat{g}_{\#}\lambda,\mathbb{P})\leqslant O\left(W_{1}^{-1/\max_{i}\{d_{i}^{C}\}}+L_{2}^{-2/\max_{i}\{d_{i}^{C}\}}+(\tau-\tau\log\tau)\right).$
@@
Here, $\widehat{g}_{i},i=1,2$ has the separable form $\left(\widehat{g}_{i}^{1}(x_{1}),\cdots,\widehat{g}_{i}^{N_{C}}(x_{N_{C}})\right)$ and $\widehat{g}_{1}^{j}\in\mathcal{NN}_{1,d_{j}^{C}}(W_{1},\Theta(d_{j}^{C}))$, $\ \widehat{g}_{2}^{j}\in\mathcal{NN}_{d_{j}^{C},d}(\Theta(d\cdot d_{j}^{C}),L_{2}),\ j\in[N_{C}]$. $\{d_{j}^{C}\}$ are the dimension of cubes in 3.4. $\widehat{g}_{3}^{\tau}$ is the Gumbel-Softmax layer with temperature parameter $\tau>0$.
Note that $\widehat{g}^{\tau}_{3}$ (the Gumbel-softmax layer) is actually a random function since the coefficient vector $\widehat{X}^{\tau}$ is a random variable. In this sense, $\widehat{g}_{\#}\lambda$ can be viewed as pushing forward uniform and Gumbel variables using a neural net.